Nutrition question, please finish it

Calculate the ATP equivalents from complete oxidat
Answer #1

Calculate the ATP equivalens from complete oxidation of 15 moles of 20:0

Number

of Carbons      Common Name            Systematic Name                      Symbol                                 Structure

20                      Arachidic acid             Eicosanoic acid                      20:0                                 

CH3(CH2)18COOH

Step 1.-

Number of Carbons/2 = Number of Acetyl CoA formed.\

=20/2 =10 Acetyl CoA

Step 2.:

= [(n/2)-1]

== [(10/2)-1] =9 rounds

Step 3 (Option a):

  (9 x 4) + (10 x 10)

step 4 = -2-----------------activation of fatty acid

Total: 134 ATP-------------------for one mole

15 moles x 134 ATP= 2010 ATP -------ANSWER